### v1

# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce a diagnostic task: file path identification from issue descriptions alone, to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions—without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
1 Introduction
As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly Srivastava et al. (2023). Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks Zhuo et al. (2024); Pan et al. (2024); Bowman & Dahl (2021). Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues Jimenez et al. (2024); OpenAI (2024). Recent advances have shown LLMs, e.g. Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress Anthropic (2025).
However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data? This concern is particularly acute because training corpora for major LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed. For instance, GitHub repositories are prominently featured in training datasets for LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2022), and Codex Chen et al. (2021), creating potential overlap between training and evaluation data.
To investigate this question systematically, we focus on a critical component of software engineering tasks: bug localization—the ability to identify which files contain bugs based on issue descriptions. This subtask is fundamental to successful issue resolution Hossain et al. (2024), yet should require genuine understanding of both the problem description and codebase structure. If models can accurately identify buggy file locations using only issue descriptions (without any repository context), this would suggest memorization of issue-file associations rather than reasoning ability.
We hypothesize that there are two types of memorization that could compromise benchmark validity: instance-specific memorization, where models recall exact issue-solution pairs from training data, and repository-bias memorization, where models have unequal familiarity with different repositories due to training data imbalances. The first type directly invalidates benchmark results by enabling models to retrieve known solutions rather than demonstrate problem-solving ability. The second type is subtler but equally problematic: while repository knowledge is valuable for LLMs, unequal repository familiarity creates systematic evaluation bias. If benchmarks like SWE-Bench predominantly sample from repositories that models know exceptionally well, they may overestimate general coding proficiency. Success might reflect disproportionate familiarity with specific codebases rather than transferable software engineering capabilities.
To systematically investigate these different memorization patterns, we design controlled experiments where models attempt to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task. Critically, performance drops by about 10% on tasks from the same repositories but not included in the SWE-Bench Verified, suggesting potential data contamination or benchmark-specific memorization rather than general coding proficiency.
To further investigate these patterns, we conduct cross-benchmark analysis comparing model performance on SWE-Bench against tasks from external repositories and different programming languages. Across ten evaluated LLMs, we consistently observe superior performance on SWE-Bench compared to external repositories, with particularly dramatic drops (up to 47 percentage points) when evaluating on repositories using different programming languages like C#. These findings raise serious concerns about the validity of current benchmark evaluations in software engineering and highlight the need for more robust, contamination-resistant evaluation approaches.
Our contributions are:
  1. 1.
    Diagnostic Task Design: We introduce a controlled bug localization task that isolates memorization effects from genuine problem-solving by asking LLMs to identify buggy files without providing repository context.
  2. 2.
    Empirical Evidence: We demonstrate through systematic experiments that current SWE-Bench performance may be inflated by memorization, with models achieving suspiciously high accuracy on context-free bug localization tasks.
  3. 3.
    Cross-Benchmark Analysis: We design a multi-benchmark validation process and provide evidence that performance disparities between SWE-Bench and external repositories suggesting repository-bias memorization rather than generalizable coding abilities, calling attention to evaluation validity concerns in the software engineering domain.
2 Approach
2.1 File Path Identification
To investigate potential memorization in large language models, we isolate a focused subtask: identifying buggy file paths from natural language issue descriptions. In our formulation, models are presented with issue reports from various software engineering benchmarks and asked to predict the file path that would require modification to resolve the described issue. Importantly, models do not receive access to repository structure, code, or metadata, only the issue text and repository name are provided. The model’s prediction is considered correct if it exactly matches one of the file paths in the ground truth (GT) patch. By narrowing down the scope to file identification and remove the repository context provided to the LLMs, we aim to expose memorization artifacts apart from their problem solving abilities. In addition, we recognize that issue descriptions themselves may contain file paths. To control for this, we also report a filtered accuracy metric on instances where such explicit path mentions are not observed, details can be found in Section 3.3.
Figure 1 illustrates the prompt template used for this task. Each prompt includes: (1) the name of the repository and (2) the issue description. The model is instructed to predict a single file path that would need to be modified to resolve the issue. Additionally, the model is asked to provide its reasoning, which enables us to analyze its underlying thought process. We show the prompt template in Figure 1 (an example of SWE-Bench instance is shown in Figure 7). We also show an example demonstrating the evaluation process in Figure 2. Here, there is only one GT path and the predicted path overlaps with the GT paths, and thus is considered correct.
You will be provided with an issue statement explaining a problem to resolve within a codebase. The code base is: scikit-learn/scikit-learn.
¡issue¿
Issue description here
¡/issue¿
DISCUSSION
Discuss here with yourself about how you came up with this response.
RESPONSE
path/to/buggy/file.py
File Path Identification Prompt Template
Figure 1: The prompt template used to test models’ ability to identify buggy files without repository access.
Ground Truth Paths: [sklearn/linear_model/ridge.py] Predicted Paths: [sklearn/linear_model/ridge.py] Is_correct: True Evaluation Demonstration
Figure 2: Example to show the evaluation process.
3 Experiment Setup
Here, we introduce the benchmarks we used for our experiment, the metrics we used to evaluate models’ performance on the path identification task, and the settings we used for the LLMs.
3.1 Benchmarks
We include three benchmarks in our preliminary experiments: SWE-Bench-Verified, SWE-Bench-C#, and RefactorBench, which we introduce below.
3.1.1 SWE-Bench-Verified
SWE-Bench-Verified is a human-validated subset of the original SWE-Bench dataset, consisting of 500 curated samples from 12 open-source Python repositories. Each task presents a GitHub issue description and requires generating code edits to resolve it. In standard evaluation, models receive both the repository snapshot and issue description. However, in our path identification experiments, we provide only the issue description and evaluate whether models can correctly identify the file paths that need modification, which should require understanding the repository structure and codebase.
Figure 3: Benchmarks’ repository distributions.
Benchmarks # of Files Issue Length
SWE-Bench Verified 763.5 451.2
SWE-Bench C# 716.7 586.2
RefactorBench 1149.2 14.6
Table 1: Average number of files and average issue description length (in tokens) per instance for each benchmark.
3.1.2 SWE-Bench-C#
SWE-Bench-C# follows the SWE-Bench construction pipeline and comprises 77 tasks from 11 C#-dominant repositories. Each task has the same format as SWE-Bench, requiring code edits based on a given issue description and repository snapshot. For our experiment, we use a subset of 40 tasks from the full dataset, and show the repository distribution in Figure 3 (a). This benchmark allows us to test whether performance advantages observed on Python-based SWE-Bench extend to different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities.
3.1.3 RefactorBench
RefactorBench Gautam et al. (2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from nine diverse open-source Python repositories. Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor of the repository. For our experiment, we use a subset of 39 tasks from the full dataset, and show the repository distribution in Figure 3 (b). Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repositories and task types. If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task-specific optimization or repository-specific optimization rather than general coding proficiency.
In addition, to ensure fair comparison across benchmarks and verify that performance differences reflect contamination rather than varying task difficulty, we analyze two key factors that could influence path identification complexity. In Table 1, we compute the average number of files with target extensions (.py for Python repositories, .cs for C# repositories) as a proxy for search space complexity, and the average issue description length as a measure of available contextual information. As shown in Table 1, SWE-Bench-Verified and SWE-Bench-C# show similar repository sizes (763.5 vs 716.7 files) and comparable issue lengths (451.2 vs 586.2 tokens), suggesting similar task complexity. On the contrary, RefactorBench’s repositories are larger (1149.2 files), and the issue descriptions are dramatically shorter (14.6 tokens), creating a more challenging scenario with larger search spaces but minimal contextual information.
3.2 Additional experiment settings
To more systematically evaluate our benchmark contamination hypothesis, we design a set of complementary test conditions beyond the SWE-Bench-Verified subset. We consider three extensions: (1) randomly sampled tasks from the SWE-Bench-Full dataset, (2) tasks from repositories not present in SWE-Bench but likely included in model training data, and (3) recently created issues from SWE-Bench repositories, which contains instances that postdate SWE-Bench’s construction. These experimental variations enable us to assess both benchmark-specific and temporal memorization patterns.
3.2.1 Full-SWE-Bench
Apart from SWE-Bench-Verified, we additionally evaluate on 200 randomly sampled instances from the full SWE-Bench Jimenez et al. (2024) dataset that are not included in the Verified subset. Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these unverified instances, it may indicate that models have been optimized specifically for the curated subset or have encountered similar patterns during training, i.e., instance-specific memorization.
3.2.2 Same-Repo Fresh Tasks
As a complementary experiment, we also collected the most recent issues from SWE-Bench repositories that were created after the original SWE-Bench dataset’s cutoff date. This includes tasks from django/django  [^1] https://github.com/django/django (183 tasks), sympy/sympy  [^2] https://github.com/sympy/sympy (119 tasks), and scikit-learn/scikit-learn  [^3] https://github.com/scikit-learn/scikit-learn(42 tasks). These fresh tasks from the same repositories in SWE-Bench provide a critical test: if models perform significantly worse on recent issues from the same repositories where they excel on SWE-Bench tasks, this would strongly suggest that their success stems from memorizing specific historical issues, instance-specific memorization, rather than repository-bias memorization.
3.2.3 External-Repo Tasks
To further test our contamination hypothesis, we constructed tasks from repositories outside the SWE-Bench dataset but with high likelihood of training data inclusion. We collected the earliest available issues and pull requests from two popular repositories: pandas-dev/pandas  [^4] https://github.com/pandas-dev/pandas (123 tasks) and pytorch/pytorch  [^5] https://github.com/pytorch/pytorch (32 tasks). These repositories are widely used, extensively documented, and have been publicly available for years, making them prime candidates for inclusion in large-scale training datasets. This setting helps us recognize repository-bias memorization. If there is strong performance on these extra-repo tasks, this would counter the possibility that models are memorizing the instances on SWE-Bench Verified. However, if models to perform significantly better on SWE-Bench tasks than on these equally-exposed repository tasks, this would indicate that that models are biased toward tasks sourced from repositories included in SWE-Bench, rather than developing general coding capabilities.
3.3 Metrics
We report accuracy as the primary metric to assess the model’s memorization of the task and its familiarity with the corresponding repository.
Accuracy (Acc.)
The accuracy is computed as the percentage of predictions that exactly match the GT patch out of the total number of samples in that category. An exact match is counted when the model-generated file path is identical to at least one of the reference GT file path(s).
Filtered accuracy (F-Acc.)
To address concerns that models might succeed through superficial pattern matching rather than genuine understanding, we introduce filtered accuracy as a control metric. This measures model performance exclusively on issue instances where problem descriptions do not explicitly contain file paths or import statements, identified through simple heuristics. By removing cases where path-like strings appear in issue descriptions, filtered accuracy isolates performance that depends purely on repository knowledge and problem-solving reasoning rather than surface-level string matching. This metric directly addresses our research question of whether models demonstrate genuine coding understanding or rely on memorized patterns.
[…]
3.4 Models
We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task. From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), the reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor OpenAI (2025a). From Anthropic, our suite comprised the Claude family: Claude 3.5 Sonnet (June 20, 2024), Claude 3.7 Sonnet (February 24, 2025), Claude 4.0 Sonnet, and Claude 4.0 Opus, both officially released on May 22, 2025.
For consistency, OpenAI’s chat models and Claude models were configured with a maximum generation length of 2,048 completion tokens. For the reasoning-centric “o” family from OpenAI (i.e., o3, o3‑mini, o4‑mini), we increased the limit to 4,096 tokens. All experiments used the respective provider APIs with default sampling settings to reflect typical usage.
4 Results
4.1 File Path Identification Accuracy
Figure 4: Comparison of model accuracy on the file path identification task across four benchmarks.
Figure 4 presents the file path identification accuracy across all models and benchmarks, revealing a clear hierarchical pattern of performance. These results provide strong evidence for the two types of memorization, instance-specific and repository-bias, hypothesized in our introduction. We analyze the evidence for each pattern below.
4.1.1 Instance-Specific memorization
The first evidence pattern emerges from performance differences within the SWE-Bench repository ecosystem. Models consistently achieve their highest accuracy on the curated SWE-Bench-Verified subset (60-76%), with a noticeable decline on the broader Full-SWE-Bench set (57-71%) and a further drop on newly collected SWE-Repo tasks (50-68%).
This graduated performance decay suggests that models’ knowledge is not uniform even across the same codebases. The superior performance on the ”Verified” set, which is the most commonly used for evaluation and reporting, points toward instance-specific memorization, where models have been disproportionately exposed to or optimized for these specific, widely-circulated benchmark problems.
4.1.2 Repository-Bias memorization
The second, more dramatic pattern appears when comparing performance between SWE-Bench and external benchmarks. While this performance drop is partly compounded by confounding variables, such as the significantly shorter issue descriptions in RefactorBench (Table 1), the data still points overwhelmingly to repository-level bias.
The most compelling evidence comes from the Outside-Repo Tasks. These tasks were sourced from popular repositories (pandas, PyTorch) that are almost certainly well-represented in training data. If models developed a general capability for popular Python projects, we would expect strong performance here. Instead, accuracy is uniformly lower (below 53%) than those in SWE-Bench Verified.
The stark contrast between high performance on the SWE-Bench repositories and poor performance on these other popular repositories refutes the idea that models are simply good at ”popular projects.” It strongly supports our hypothesis of repository-bias memorization, where models have overfit to the specific architectural patterns and problem distributions of the dozen repositories in the SWE-Bench collection, failing to develop transferable skills.
4.1.3 Model-Specific Contamination Patterns
The results reveal differences in contamination patterns across model families and vendors:
OpenAI chat Models: OpenAI’s chat model variants show inconsistent performance patterns across benchmarks, with no stable ranking order. The performance ranking of these models varies significantly between different benchmark types. The only consistent pattern is GPT-4o-2024-08-06, which maintains mediocre performance across most evaluations.
OpenAI reasoning Models: While o3 demonstrates higher accuracy than o3-mini and o4-mini due to its enhanced capabilities, the performance patterns of o3-mini and o4-mini follow similar trends to the chat models. We cannot identify a consistently stronger model within this family for the path prediction task across different benchmarks.
Anthropic Models: Claude models exhibit more consistent performance patterns, with newer model generations generally showing higher path prediction accuracy. However, Claude Sonnet 4 shows notably lower performance than other Claude family models on RefactorBench and Extra-Repo tasks, which represents an exception to the otherwise consistent generational improvement pattern. Overall, Claude Sonnet 4 and Opus 4 maintain relatively higher performance across most benchmarks compared to other Claude models.
Cross-Vendor Consistency: A key finding from our evaluation of ten different LLMs is the remarkable consistency of these patterns across different model families and vendors. The fact that all models, regardless of vendor, exhibit the same performance hierarchy (SWE-bench Verified >>> Full-SWE-Bench >>> SWE-Repo Tasks >>> SWE-bench C#, RefactorBench, and Outside-Repo Tasks) indicates the performance disparities reflect systematic exposure patterns in training data rather than isolated vendor-specific issues.
4.1.4 Implications for Benchmark Validity
The results provide strong empirical evidence for the two forms of memorization hypothesized in our introduction. The performance degradation within the SWE-Bench ecosystem points to instance-specific memorization, while the more significant performance collapse on external repositories indicates repository-bias memorization. The dual presence of these phenomena suggests that high scores on SWE-Bench are not purely indicative of generalizable coding ability. Instead, they are likely inflated by these confounding factors, reflecting a specialized knowledge of the benchmark’s specific instances and repository patterns rather than transferable software engineering skills.
Dataset Mentioned # (%) Unmentioned # (%) Total #
SWE-bench Verified 135 (27.0%) 365 (73.0%) 500
Full-SWE-Bench 42 (21.0%) 158 (79.0%) 200
SWE-Repo Tasks 100 (29.1%) 244 (70.9%) 344
SWE-bench C# 11 (14.7%) 64 (83.1%) 77
RefactorBench 1 (2.6%) 38 (97.4%) 39
Outside-Repo Tasks 26 (16.5%) 132 (83.5%) 158
Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contains ground truth path information and those do not.
4.2 File Path Identification Accuracy on Filtered Instances
To more rigorously test our contamination hypothesis, this section examines model performance only on instances where the issue description provides no explicit file paths or import statements. The results, presented in Figure 5, remove superficial string matching as a confounding variable and reveal that the underlying memorization patterns persist.
First, models still achieve their highest accuracy on SWE-Bench Verified instances, followed by Full SWE-Bench and SWE-Repo tasks confirming a pattern of instance-specific memorization that is not dependent on clues from the prompt. Second, the sharp performance drop on external repositories also persists, providing definitive evidence for repository-bias memorization.
The persistence of these patterns in the filtered data is our most critical finding. It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns. This strongly suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
Figure 5: Comparison of models’ filtered accuracy on the file path identification task across benchmarks.
5 Related Work
5.1 Coding Benchmarks and Agents
Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’ coding abilities, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval Jimenez et al. (2024); Zhuo et al. (2024); Xia et al. (2024). Among them, SWE-Bench has emerged as one of the most popular evaluation method for LLM’s coding abilities (Jimenez et al., 2024). The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting. However, the public nature of GitHub data raises concerns about potential exposure during training. Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al., 2024; Wang et al., 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
5.2 Benchmark Contamination
The issue of benchmark contamination, where models are inadvertently trained on evaluation data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding Zhou et al. (2023). To specifically combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
Task mutation approaches aim to detect memorization by altering existing benchmark problems. For example, EvoEval Xia et al. (2024) uses LLMs to create semantic-altering transformations of HumanEval Chen et al. (2021) problems, while other methods Chen et al. (2025) use AST-based mutations to measure how performance changes with structural code modifications. Similarly, TaskEval Tambon et al. (2025) measures prompt sensitivity by creating numerous variations of the same task. These methods are effective at revealing overfitting on specific problem formulations.
Metric-based probing approaches, in contrast, analyze the model’s outputs without changing the task itself. These techniques Li (2023); Ramos et al. (2025) include using perplexity analysis or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if a model has likely seen a specific bug or solution before. However, perplexity and NLL approaches cannot be applied to commercial models, for we do not have access to the model’s hidden states. N-gram similarity, while applicable, can be a noisy indicator for complex code patches; any functionally correct solution will inevitably share significant textual overlap with the ground truth, making it difficult to distinguish genuine problem-solving from verbatim memorization.
While existing mutation and metric-based methods are effective for single-function code generation, they primarily target instance-specific memorization and are less suited for complex, multi-file software engineering tasks like those in SWE-Bench. Our work introduces a complementary diagnostic approach that targets a different type of contamination: repository-bias memorization. Through subtask isolation—testing models on prerequisite tasks without necessary context—and systematic cross-benchmark analysis across different repository sets and programming languages, we provide a methodology to detect whether high benchmark performance reflects genuine problem-solving ability or disproportionate familiarity with specific codebases.
6 Conclusion
This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem, and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points) despite their popularity and likely training data inclusion. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
These findings suggest that widely-cited improvements in SWE-Bench performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities. The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills. Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than dataset-specific artifacts.
Appendix A Appendix
A.1 Full Prompt Template
def get_file_path_prompt(entry):
    file_prompt = ’’’You will be provided with an issue statement explaining a problem to resolve within a codebase. The code base is: {0}.
   <issue>
   {1}
   </issue>
   You must always include one discussion and one response as part of your response.
   Make sure you do not have multiple discussion/response tags.
   Please make sure your output precisely matches the following format.
   DISCUSSION
   Discuss here with yourself about how you came up with this response.
   RESPONSE
   response(s) to the question asked.
   Now answer the following question:
   Given the issue description and project, provide a file-path of the .py file containing the issue relative to the root.’’’.format(entry[’repo’], entry[’problem_statement’])
    model_input = [{
       ’role’: ’user’,
       ’content’: file_prompt
    return model_input
Complete File Path Prompt Implementation
Figure 6: The complete implementation of the file path identification prompt used in our experiments.
RidgeClassifierCV store_cv_values Issue Error: TypeError: __init__() got an unexpected keyword argument ’store_cv_values’ Repro: from sklearn import linear_model import numpy as np x = np.random.randn(100, 30) y = np.random.normal(size=100) model = linear_model.RidgeClassifierCV( alphas=np.arange(0.1, 1000, 0.1), normalize=True, store_cv_values=True ).fit(x, y) Note: store_cv_values is not a valid parameter, yet some attributes depend on it.
Figure 7: Minimal reproduction of RidgeClassifierCV parameter issue
A.2 Additional Results
[Additional tables and figures supporting the main findings]
Model Mentioned Not Mentioned Overall
Single GT Multi GT Overall Single GT Multi GT Overall Single GT Multi GT Overall
claude-3-5-sonnet 88.60 71.43 85.93 51.90 44.90 50.96 61.63 52.86 60.40
claude-3-7-sonnet 92.11 85.71 91.11 60.13 57.14 59.73 68.60 65.71 68.20
claude-sonnet-4 91.23 85.71 90.37 65.51 53.06 63.84 72.33 62.86 71.00
claude-opus-4 92.11 95.24 92.59 67.41 61.22 66.58 73.95 71.43 73.60
gpt-4o (May) 91.23 85.71 90.37 51.27 55.10 51.78 61.86 64.29 62.20
gpt-4o 88.60 80.95 87.41 53.16 51.02 52.88 62.56 60.00 62.20
gpt-4.1 85.09 71.43 82.96 56.65 55.10 56.44 64.19 60.00 63.60
o3-mini 91.23 85.71 90.37 60.13 63.27 60.55 68.37 70.00 68.60
o3 92.98 76.19 90.37 71.84 65.31 70.96 77.44 68.57 76.20
o4-mini 92.98 80.95 91.11 59.49 48.98 58.08 68.37 58.57 67.00
Table 3: Accuracy of model-predicted file paths on SWE-Bench Verified, stratified by whether any GT path is explicitly mentioned in the problem description. Results are further broken down into three categories: Single GT Path (tasks where the GT patch modifies only a single execution path), Multiple GT Paths (tasks involving modifications across multiple execution paths), and Overall (all tasks). Among the task instances, 135 mention at least one GT path, while 365 do not.

### v2

@@
 It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts.
-In this work, we introduce a diagnostic task: file path identification from issue descriptions alone, to probe models’ underlying knowledge.
+In this work, we introduce two diagnostic task: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge.
 We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving.
-We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions—without access to repository structure.
+We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure.
 This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization.
+Similar patterns is also observed for function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks.
 These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
@@
 (2021), creating potential overlap between training and evaluation data.
-To investigate this question systematically, we focus on a critical component of software engineering tasks: bug localization—the ability to identify which files contain bugs based on issue descriptions.
-This subtask is fundamental to successful issue resolution Hossain et al.
-(2024), yet should require genuine understanding of both the problem description and codebase structure.
-If models can accurately identify buggy file locations using only issue descriptions (without any repository context), this would suggest memorization of issue-file associations rather than reasoning ability.
+To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs based on issue descriptions, and patch generation, the ability to correct the buggy code.
+Both subtasks are fundamental to successful issue resolution.
+For bug localization Hossain et al.
+(2024), models should require genuine understanding of both the problem description and codebase structure to identify relevant files.
+For patch generation, models need to understand the semantics of the code and, potentially with additional context from the issue description, reason to generate the correct solution.
+Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability.
+However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns and domain knowledge during training.
+To address this challenge, we propose a differential testing approach inspired by software engineering practices: when ground truth oracles are unavailable, we use comparative analysis across similar systems as proxies.
+In our case, we evaluate models across multiple similar benchmarks and repositories, using performance disparities as indicators of memorization.
+Genuine coding ability should produce relatively consistent performance across comparable tasks, while memorization effects would manifest as systematic performance advantages on specific benchmarks where training data overlap is more likely.
 We hypothesize that there are two types of memorization that could compromise benchmark validity: instance-specific memorization, where models recall exact issue-solution pairs from training data, and repository-bias memorization, where models have unequal familiarity with different repositories due to training data imbalances.
@@
 Success might reflect disproportionate familiarity with specific codebases rather than transferable software engineering capabilities.
-To systematically investigate these different memorization patterns, we design controlled experiments where models attempt to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context.
-Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task.
-Critically, performance drops by about 10% on tasks from the same repositories but not included in the SWE-Bench Verified, suggesting potential data contamination or benchmark-specific memorization rather than general coding proficiency.
-To further investigate these patterns, we conduct cross-benchmark analysis comparing model performance on SWE-Bench against tasks from external repositories and different programming languages.
-Across ten evaluated LLMs, we consistently observe superior performance on SWE-Bench compared to external repositories, with particularly dramatic drops (up to 47 percentage points) when evaluating on repositories using different programming languages like C#.
+To systematically investigate these different memorization patterns, we design a controlled experiment pipeline that includes two diagnostic tasks: file-path identification and function reproduction.
+The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context.
+The function reproduction task measures verbatim similarity between generated patches and ground truth solutions.
+Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances for file-path identification, despite lacking the contextual information that should be necessary for this task.
+More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
 These findings raise serious concerns about the validity of current benchmark evaluations in software engineering and highlight the need for more robust, contamination-resistant evaluation approaches.
@@
   1. 1.
-    Diagnostic Task Design: We introduce a controlled bug localization task that isolates memorization effects from genuine problem-solving by asking LLMs to identify buggy files without providing repository context.
+    Differential Testing for Memorization Detection: We introduce a systematic approach using performance disparities across similar benchmarks as memorization proxies when direct measurement is impossible.
   2. 2.
-    Empirical Evidence: We demonstrate through systematic experiments that current SWE-Bench performance may be inflated by memorization, with models achieving suspiciously high accuracy on context-free bug localization tasks.
+    Evidence of Benchmark Compromise: We demonstrate that state-of-the-art models achieve suspiciously high performance on SWE-Bench (76% context-free accuracy, elevated 5-gram similarity) with substantial drops on external benchmarks, indicating memorization rather than genuine capability.
   3. 3.
-    Cross-Benchmark Analysis: We design a multi-benchmark validation process and provide evidence that performance disparities between SWE-Bench and external repositories suggesting repository-bias memorization rather than generalizable coding abilities, calling attention to evaluation validity concerns in the software engineering domain.
+    Automated Evaluation Framework: We provide a reusable protocol for detecting memorization in any coding benchmark by comparing performance across similar evaluation sets.
 2 Approach
@@
 Figure 2: Example to show the evaluation process.
+2.2 Function Reproduction Task
+For the function reproduction task, we filter instances to include only function-level modifications and present models with issue descriptions alongside pre-patched files where modified functions are completely removed.
+Models must generate complete implementations for each missing function.
+Since function specification is unavailable, successful reproduction indicates memorization of specific textual sequences rather than contextual reasoning.
+We compare performance across multiple benchmarks to isolate memorization effects from general programming ability.
+We show the prompt template we used for function reproduction task in Appendix (Figure9).
+Our approach directly tests memorization by requiring models to reproduce complete function implementations without access to signatures, docstrings, or relevant context from other files.
+This constraint ensures that successful reproduction requires prior exposure to the specific implementation rather than general problem-solving ability.
+The high consecutive 5-gram accuracy on SWE-Bench Verified compared to functionally equivalent control tasks provides evidence of memorized textual patterns rather than emergent programming competence.
 3 Experiment Setup
@@
 3.3 Metrics
-We report accuracy as the primary metric to assess the model’s memorization of the task and its familiarity with the corresponding repository.
+In this section, we details our choice of metrics used for our experiments.
 Accuracy (Acc.)
@@
 […]
+5-gram Consecutive Accuracy (5-gram Acc.)
+To quantify the degree of verbatim memorization in model outputs, we compute 5-gram consecutive accuracy, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the ground truth.
+This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching to prevent inflation from repeated phrases.
+Given that random 5-gram matches are statistically improbable, high scores on this metric provide strong evidence of memorization rather than genuine code generation.
+This allows us to distinguish models that have memorized data patterns in the training samples.
 3.4 Models
@@
 Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contains ground truth path information and those do not.
+Figure 5: Comparison of models’ filtered accuracy on the file path identification task across benchmarks.
 4.2 File Path Identification Accuracy on Filtered Instances
@@
 It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns.
-This strongly suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
-Figure 5: Comparison of models’ filtered accuracy on the file path identification task across benchmarks.
+This suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
+Figure 6: Comparison of models’ consecutive 5-gram accuracy on function reproduction across benchmarks.
+Red ticks indicate improved verbatim similarity with the ground truth code, while blue ticks indicate a decrease.
+4.3 Function Reproduction Result
+We analyze model performance on the function reproduction task in this subsection.
+As shown in Figure 6, models achieve up to 35.0% consecutive 5-gram accuracy reproducing ground truth functions (the lower plot) from SWE-Bench Verified compared to up to 18.2% on instances outside the SWE-Bench ecosystem666 Due to limited resources, we only conducted experiments for o3-mini for the OpenAI reasoning models.
+.
+With limited context provided, only function names and issue descriptions, models lack sufficient information to systematically derive correct implementations.
+High consecutive 5-gram accuracy under these constraints requires reproducing exact textual sequences, indicating reliance on memorized training data rather than reasoning.
+Instance-Specific Memorization
+: Performance reveals a clear hierarchy that mirrors benchmark curation processes: SWE Verified (35.0%) >>> SWE Full (28.9%) >>> SWE Extra (18.3%).
+Critically, SWE Extra performance drops to the same level as external benchmarks (RefactorBench: 18.2%, Outside Tasks: 17.7%), despite originating from identical repositories as SWE Verified.
+This pattern indicates that models have specifically memorized the curated, canonical solutions that human evaluators deemed high-quality, rather than developing general familiarity with repository codebases.
+These results demonstrate that SWE-Bench Verified performance substantially reflects memorization of training sequences rather than coding competence.
+The inability to reproduce functions from less-curated instances within the same repositories, combined with performance equivalence between SWE Extra and external benchmarks, reveals that apparent programming ability is largely an artifact of exposure to the specific, high-quality solutions selected for benchmark inclusion.
 5 Related Work
@@
 N-gram similarity, while applicable, can be a noisy indicator for complex code patches; any functionally correct solution will inevitably share significant textual overlap with the ground truth, making it difficult to distinguish genuine problem-solving from verbatim memorization.
-While existing mutation and metric-based methods are effective for single-function code generation, they primarily target instance-specific memorization and are less suited for complex, multi-file software engineering tasks like those in SWE-Bench.
-Our work introduces a complementary diagnostic approach that targets a different type of contamination: repository-bias memorization.
-Through subtask isolation—testing models on prerequisite tasks without necessary context—and systematic cross-benchmark analysis across different repository sets and programming languages, we provide a methodology to detect whether high benchmark performance reflects genuine problem-solving ability or disproportionate familiarity with specific codebases.
+While existing mutation and metric-based methods are effective for single-function code generation, they face limitations when applied to complex, multi-file software engineering benchmarks like SWE-Bench.
+Task mutation approaches require generating semantically equivalent variations of real-world GitHub issues, a significantly more challenging task than mutating isolated coding problems.
+Metric-based approaches using perplexity or NLL cannot be applied to commercial models due to lack of access to hidden states, while n-gram similarity becomes noisy for complex patches where any functionally correct solution will share substantial textual overlap with ground truth.
+Our work introduces a novel differential testing framework that addresses these limitations by using comparative performance analysis across related benchmarks as memorization proxies.
+Rather than requiring ground truth knowledge of training data exposure, we systematically compare model performance across benchmark variants and external evaluations to detect suspicious performance patterns.
+This differential testing methodology provides a systematic, reusable framework for detecting memorization in any coding benchmark without requiring access to training data or model internals.
 6 Conclusion
 This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization.
-Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem, and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points) despite their popularity and likely training data inclusion.
+Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem (in both file-path identification task and function reproduction task), and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points for file-path identification task) despite their popularity and likely training data inclusion.
 The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
-These findings suggest that widely-cited improvements in SWE-Bench performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities.
+These findings suggest that the improvements in SWE-Bench Verified performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities.
 The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills.
@@
 Among the task instances, 135 mention at least one GT path, while 365 do not.
+You are provided with a code repository and an issue description.
+Your task is to implement the complete function bodies for the marked RESPONSE comments.
+These functions are necessary to resolve the issue described in the problem statement.
+Repository: scikit-learn/scikit-learn <issue> Issue description here </issue> ...
+HINT: Fixes in _tags.py: modification in get_tags === File: sklearn/utils/_tags.py === <file> ...
+#TODO: RESPONSE 1: ...
+</file> ...
+You must provide complete implementations for each RESPONSE marker.
+Each function should: 1.
+Be properly indented to match its position in the file 2.
+Include the complete function body including the function signature (def line) 3.
+Address the issue described above RESPONSE INSTRUCTIONS: RESPONSE 1: Complete implementation of get_tags in sklearn/utils/_tags.py IMPORTANT: - Provide the COMPLETE function implementation, including the ’def’ line and all decorators if any - Maintain correct Python indentation - Each RESPONSE should contain the entire function from decorators (if any) through the complete function body - Number your responses to match the RESPONSE markers in the code - Wrap each response in triple backticks (‘‘‘) Please think through the issue first, then provide your implementations.
+DISCUSSION: Analyze the issue and plan your implementations here RESPONSE 1: ‘‘‘python # Your implementation here ‘‘‘ Function Reproduction Task Prompt Template
+Figure 9: The prompt template used to test models’ ability to reproduce ground truth function without function specification.

### v3

@@
 For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues.
-Recent LLMs show impressive performance on SWE-Bench, leading to optimism about their capacity for complex coding tasks.
+Recent LLMs show impressive performance on SWE-Bench Verified, leading to optimism about their capacity for complex coding tasks.
 However, current evaluation protocols may overstate these models’ true capabilities.
 It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts.
-In this work, we introduce two diagnostic task: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge.
-We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving.
-We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure.
+In this work, we introduce two new diagnostic tasks: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge.
+We present empirical evidence that performance gains on SWE-Bench Verified may be partially driven by memorization rather than genuine problem-solving.
+We show that state-of-the-art (SoTA) models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure.
 This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization.
-Similar patterns is also observed for function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks.
+Similar patterns are also observed for the function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks (up to 35% consecutive 5-gram overlap ratio on SWE-Bench Verified and Full, but only up to 18% for tasks in other benchmarks).
 These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
 1 Introduction
-As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly Srivastava et al.
-(2023).
-Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks Zhuo et al.
-(2024); Pan et al.
-(2024); Bowman & Dahl (2021).
-Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues Jimenez et al.
-(2024); OpenAI (2024).
+As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly (Srivastava et al.
+2023).
+Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks (Zhuo et al.
+2024; Pan et al.
+2024; Bowman and Dahl 2021).
+Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues (Jimenez et al.
+2024; OpenAI 2024).
 Recent advances have shown LLMs, e.g.
-Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress Anthropic (2025).
+Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress (Anthropic 2025).
 However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data?
-This concern is particularly acute because training corpora for major LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed.
-For instance, GitHub repositories are prominently featured in training datasets for LLaMA Touvron et al.
-(2023), PaLM Chowdhery et al.
-(2022), and Codex Chen et al.
-(2021), creating potential overlap between training and evaluation data.
-To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs based on issue descriptions, and patch generation, the ability to correct the buggy code.
-Both subtasks are fundamental to successful issue resolution.
-For bug localization Hossain et al.
-(2024), models should require genuine understanding of both the problem description and codebase structure to identify relevant files.
-For patch generation, models need to understand the semantics of the code and, potentially with additional context from the issue description, reason to generate the correct solution.
+This concern is particularly acute because training corpora for LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed.
+For instance, GitHub repositories are prominently featured in training datasets for LLaMA (Touvron et al.
+2023), PaLM (Chowdhery et al.
+2022), and Codex (Chen et al.
+2021), creating potential overlap between training and evaluation data.
+To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs, and patch generation, the ability to generate the corrected version of the buggy code.
+Both sub-tasks are fundamental to successful issue resolution.
+For bug localization (Hossain et al.
+2024), models should require a genuine understanding of both the problem description and codebase structure to identify relevant files.
+For patch generation, models need to understand the semantics of the code and, based on the issue description to reason and generate the correct solution.
 Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability.
-However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns and domain knowledge during training.
-To address this challenge, we propose a differential testing approach inspired by software engineering practices: when ground truth oracles are unavailable, we use comparative analysis across similar systems as proxies.
-In our case, we evaluate models across multiple similar benchmarks and repositories, using performance disparities as indicators of memorization.
-Genuine coding ability should produce relatively consistent performance across comparable tasks, while memorization effects would manifest as systematic performance advantages on specific benchmarks where training data overlap is more likely.
-We hypothesize that there are two types of memorization that could compromise benchmark validity: instance-specific memorization, where models recall exact issue-solution pairs from training data, and repository-bias memorization, where models have unequal familiarity with different repositories due to training data imbalances.
-The first type directly invalidates benchmark results by enabling models to retrieve known solutions rather than demonstrate problem-solving ability.
-The second type is subtler but equally problematic: while repository knowledge is valuable for LLMs, unequal repository familiarity creates systematic evaluation bias.
-If benchmarks like SWE-Bench predominantly sample from repositories that models know exceptionally well, they may overestimate general coding proficiency.
-Success might reflect disproportionate familiarity with specific codebases rather than transferable software engineering capabilities.
-To systematically investigate these different memorization patterns, we design a controlled experiment pipeline that includes two diagnostic tasks: file-path identification and function reproduction.
+To systematically investigate these different memorization patterns, we design three diagnostic tasks: file-path identification, function reproduction, and prefix completion.
 The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context.
-The function reproduction task measures verbatim similarity between generated patches and ground truth solutions.
-Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances for file-path identification, despite lacking the contextual information that should be necessary for this task.
+The function reproduction task measures models’ ability to reproduce fixed function code without providing the necessary information for such reproduction, such as the respective function specification.
+The prefix completion task checks models’ ability to verbatim generate the solutions of tasks given the tasks’ prefixes.
+However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns during training, and there lack of a standard baseline for their understanding of general coding abilities.
+To address this challenge, we propose a cross-benchmark performance analysis approach and use comparative analysis across similar systems as proxies.
+In our case, we evaluate models across multiple similar benchmarks and repositories, using performance difference across similar tasks as indicators of memorization.
+Genuine coding skill should lead to similar performance across comparable tasks, while memorization would result in unusually high scores on tasks that likely overlap with the model’s training data.
+Our results reveal concerning patterns: on the file-path identification task, SoTA models like o3 (OpenAI 2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task.
 More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
 These findings raise serious concerns about the validity of current benchmark evaluations in software engineering and highlight the need for more robust, contamination-resistant evaluation approaches.
-Our contributions are:
+Our contributions includes:
   1. 1.
-    Differential Testing for Memorization Detection: We introduce a systematic approach using performance disparities across similar benchmarks as memorization proxies when direct measurement is impossible.
+    Cross-Benchmark Performance Analysis for Memorization Detection: We introduce a systematic approach for benchmark memorization detection using performance disparities across similar benchmarks.
+When models exhibit disproportionately high performance on specific benchmarks compared to similar tasks, this indicates potential training data contamination rather than genuine coding ability.
   2. 2.
-    Evidence of Benchmark Compromise: We demonstrate that state-of-the-art models achieve suspiciously high performance on SWE-Bench (76% context-free accuracy, elevated 5-gram similarity) with substantial drops on external benchmarks, indicating memorization rather than genuine capability.
+    Diagnostic Task Design: We introduce two controlled diagnostic tasks: file-path identification without repository context and verbatim patch reproduction analysis, that isolate and measure different types of memorization effects in software engineering benchmarks.
   3. 3.
-    Automated Evaluation Framework: We provide a reusable protocol for detecting memorization in any coding benchmark by comparing performance across similar evaluation sets.
+    Evidence of Benchmark Compromise: We show that models achieve suspiciously high performance on SWE-Bench (76% context-free accuracy, elevated 5-gram score) with substantial drop on external benchmarks, indicating memorization rather than genuine capability.
 2 Approach
-2.1 File Path Identification
-To investigate potential memorization in large language models, we isolate a focused subtask: identifying buggy file paths from natural language issue descriptions.
-In our formulation, models are presented with issue reports from various software engineering benchmarks and asked to predict the file path that would require modification to resolve the described issue.
-Importantly, models do not receive access to repository structure, code, or metadata, only the issue text and repository name are provided.
-The model’s prediction is considered correct if it exactly matches one of the file paths in the ground truth (GT) patch.
-By narrowing down the scope to file identification and remove the repository context provided to the LLMs, we aim to expose memorization artifacts apart from their problem solving abilities.
-In addition, we recognize that issue descriptions themselves may contain file paths.
-To control for this, we also report a filtered accuracy metric on instances where such explicit path mentions are not observed, details can be found in Section 3.3.
-Figure 1 illustrates the prompt template used for this task.
-Each prompt includes: (1) the name of the repository and (2) the issue description.
-The model is instructed to predict a single file path that would need to be modified to resolve the issue.
-Additionally, the model is asked to provide its reasoning, which enables us to analyze its underlying thought process.
-We show the prompt template in Figure 1 (an example of SWE-Bench instance is shown in Figure 8).
-We also show an example demonstrating the evaluation process in Figure 2.
-Here, there is only one GT path and the predicted path overlaps with the GT paths, and thus is considered correct.
+Figure 1: Overview of our benchmark memorization detection approach.
 You will be provided with an issue statement explaining a problem to resolve within a codebase. The code base is: scikit-learn/scikit-learn.
@@
 File Path Identification Prompt Template
-Figure 1: The prompt template used to test models’ ability to identify buggy files without repository access.
-Ground Truth Paths: [sklearn/linear_model/ridge.py] Predicted Paths: [sklearn/linear_model/ridge.py] Is_correct: True Evaluation Demonstration
-Figure 2: Example to show the evaluation process.
+Figure 2: The prompt template used to test models’ ability to identify buggy files without repo access.
+In this section, we introduce the three diagnostic tasks designed for the detection of memorization in software engineering benchmarks: file path identification tests whether models can locate buggy files without repo context, function reproduction checks complete implementation generation without specifications, and prefix completion directly measures verbatim code sequence reproduction.
+2.1 File Path Identification Task
+In this task, models are presented with issue reports and asked to predict the file path that would require modification to resolve the described issue.
+Importantly, models do not receive access to repo structure, code, or metadata; only the issue text and repo name are provided.
+The model’s prediction is considered correct if it exactly matches one of the file paths in the ground truth (GT) patch.
+By narrowing down the scope to file identification and removing the repo context provided to the LLMs, we aim to expose memorization artifacts apart from their problem-solving abilities.
+In addition, we recognize that issue descriptions themselves may contain file paths.
+To resolve this, we also report a filtered accuracy metric on instances where such explicit path mentions are not observed; details available in Section 3.3.
+Figure 2 illustrates the prompt template used for this task.
+Each prompt includes: (1) the name of the repo and (2) the issue description.
+The model is instructed to predict a single file path that would need to be modified to resolve the issue.
+Additionally, the model is asked to provide its reasoning, which enables us to analyze its underlying thought process.
+We show the prompt template in Figure 2.
 2.2 Function Reproduction Task
-For the function reproduction task, we filter instances to include only function-level modifications and present models with issue descriptions alongside pre-patched files where modified functions are completely removed.
+In Function Reproduction Task, we analyze whether models can generate complete functions that are modified by the GT patch without having knowledge of its specification or function signature.
+Specifically, we present models with issue descriptions alongside buggy files where functions modified by the GT patch are completely removed.
 Models must generate complete implementations for each missing function.
-Since function specification is unavailable, successful reproduction indicates memorization of specific textual sequences rather than contextual reasoning.
-We compare performance across multiple benchmarks to isolate memorization effects from general programming ability.
-We show the prompt template we used for function reproduction task in Appendix (Figure9).
+Since the function specification is unavailable, successful reproduction indicates memorization of specific textual sequences rather than problem-solving ability.
+We compare performance across benchmarks to isolate memorization effects from general programming ability.
+The prompt template is available in supplementary materials.
 Our approach directly tests memorization by requiring models to reproduce complete function implementations without access to signatures, docstrings, or relevant context from other files.
-This constraint ensures that successful reproduction requires prior exposure to the specific implementation rather than general problem-solving ability.
-The high consecutive 5-gram accuracy on SWE-Bench Verified compared to functionally equivalent control tasks provides evidence of memorized textual patterns rather than emergent programming competence.
+We note that this is a single round generation, and the models must rely only on the provided file content to reproduce the removed functions.
+The models do not have access to the repo.
+These constraints ensures that successful reproduction requires prior exposure to the specific implementation rather than general problem-solving ability.
+2.3 Prefix Completion Task
+Additionally, we designed a controlled experiment testing whether models can reproduce exact code sequences in the task solution given only the prefix code. The task proceeds in three phases:
+Prefix Extraction: For each instance in SWE-Bench Verified, we extract the lines of code before each buggy code snippet that was modified by the GT patch.
+Prompt Construction: The extracted code prefix is provided to the model, along with instructions to complete the remaining code implementation.
+Generation and Evaluation: Models generate continuations based solely on the provided prefix.
+We then compare the generated code against the GT to identify verbatim matches for up to $N$ lines of code.
+$N$ is the number of lines that are modified by the GT patch at the respective location.
+This Prefix Completion Task complements the Function Reproduction Task by testing memorization in a setting that directly mirrors the models’ auto-regressive pre-training.
+Together, these two tasks assess memorization at two different scales: the recall of complete functions versus the verbatim completion of sequential code.
 3 Experiment Setup
-Here, we introduce the benchmarks we used for our experiment, the metrics we used to evaluate models’ performance on the path identification task, and the settings we used for the LLMs.
+We introduce the benchmarks and metrics we used to evaluate the models, and the settings for model inference.
 3.1 Benchmarks
-We include three benchmarks in our preliminary experiments: SWE-Bench-Verified, SWE-Bench-C#, and RefactorBench, which we introduce below.
-3.1.1 SWE-Bench-Verified
+We include three established benchmarks: SWE-Bench-Verified, SWE-Bench-C#, and RefactorBench.
+SWE-Bench-Verified
 SWE-Bench-Verified is a human-validated subset of the original SWE-Bench dataset, consisting of 500 curated samples from 12 open-source Python repositories.
 Each task presents a GitHub issue description and requires generating code edits to resolve it.
-In standard evaluation, models receive both the repository snapshot and issue description.
-However, in our path identification experiments, we provide only the issue description and evaluate whether models can correctly identify the file paths that need modification, which should require understanding the repository structure and codebase.
-Figure 3: Benchmarks’ repository distributions.
+In standard settings, models are given access to the repo and the issue description.
+However, in our experiments, we provide only context given by the task as detailed in Section 2.
+SWE-Bench-C#
+SWE-Bench-C# is an internal benchmark that follows SWE-Bench’s construction pipeline and comprises 75 tasks from eleven C# repositories.
+Each task has the same format as SWE-Bench, requiring code edits based on a given issue description.
+This benchmark allows us to test whether performance patterns observed on Python-based tasks extend to tasks from different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities.
+Details are available in the supplementary materials.
+RefactorBench
+RefactorBench (Gautam et al.
+2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from 9 open-source Python repos.
+Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor.
+For our experiment, we use a subset of 39 tasks from the full dataset.
+Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repos and task types.
+If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task or repo-specific optimization rather than general coding proficiency.
 Benchmarks # of Files Issue Length
@@
 Table 1: Average number of files and average issue description length (in tokens) per instance for each benchmark.
-3.1.2 SWE-Bench-C#
-SWE-Bench-C# follows the SWE-Bench construction pipeline and comprises 77 tasks from 11 C#-dominant repositories.
-Each task has the same format as SWE-Bench, requiring code edits based on a given issue description and repository snapshot.
-For our experiment, we use a subset of 40 tasks from the full dataset, and show the repository distribution in Figure 3 (a).
-This benchmark allows us to test whether performance advantages observed on Python-based SWE-Bench extend to different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities.
-3.1.3 RefactorBench
-RefactorBench Gautam et al.
-(2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from nine diverse open-source Python repositories.
-Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor of the repository.
-For our experiment, we use a subset of 39 tasks from the full dataset, and show the repository distribution in Figure 3 (b).
-Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repositories and task types.
-If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task-specific optimization or repository-specific optimization rather than general coding proficiency.
-In addition, to ensure fair comparison across benchmarks and verify that performance differences reflect contamination rather than varying task difficulty, we analyze two key factors that could influence path identification complexity.
-In Table 1, we compute the average number of files with target extensions (.py for Python repositories, .cs for C# repositories) as a proxy for search space complexity, and the average issue description length as a measure of available contextual information.
-As shown in Table 1, SWE-Bench-Verified and SWE-Bench-C# show similar repository sizes (763.5 vs 716.7 files) and comparable issue lengths (451.2 vs 586.2 tokens), suggesting similar task complexity.
-On the contrary, RefactorBench’s repositories are larger (1149.2 files), and the issue descriptions are dramatically shorter (14.6 tokens), creating a more challenging scenario with larger search spaces but minimal contextual information.
+Additionally, to ensure fair comparison across benchmarks and verify that performance differences reflect contamination rather than varying task difficulty, we analyze two key factors that may influence path identification complexity.
+In Table 1, we compute the average number of files with target extensions (.py for Python, .cs for C# repositories) as a proxy for search space complexity, and average issue description length as a measure of available contextual info.
+As seen in Table 1, SWE-Bench-Verified and SWE-Bench-C# have similar repo sizes (763.5 vs 716.7 files) and comparable issue lengths (451.2 vs 586.2 tokens), suggesting similar task complexity.
+In contrast, RefactorBench repos are larger (1149.2 files), and the issues are much shorter (14.6 tokens), creating a more challenging scenario with larger search spaces but minimal contextual information.
 3.2 Additional experiment settings
-To more systematically evaluate our benchmark contamination hypothesis, we design a set of complementary test conditions beyond the SWE-Bench-Verified subset.
-We consider three extensions: (1) randomly sampled tasks from the SWE-Bench-Full dataset, (2) tasks from repositories not present in SWE-Bench but likely included in model training data, and (3) recently created issues from SWE-Bench repositories, which contains instances that postdate SWE-Bench’s construction.
-These experimental variations enable us to assess both benchmark-specific and temporal memorization patterns.
-3.2.1 Full-SWE-Bench
-Apart from SWE-Bench-Verified, we additionally evaluate on 200 randomly sampled instances from the full SWE-Bench Jimenez et al.
-(2024) dataset that are not included in the Verified subset.
-Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these unverified instances, it may indicate that models have been optimized specifically for the curated subset or have encountered similar patterns during training, i.e., instance-specific memorization.
-3.2.2 Same-Repo Fresh Tasks
-As a complementary experiment, we also collected the most recent issues from SWE-Bench repositories that were created after the original SWE-Bench dataset’s cutoff date.
-This includes tasks from django/django  [^1] https://github.com/django/django (183 tasks), sympy/sympy  [^2] https://github.com/sympy/sympy (119 tasks), and scikit-learn/scikit-learn  [^3] https://github.com/scikit-learn/scikit-learn(42 tasks).
-These fresh tasks from the same repositories in SWE-Bench provide a critical test: if models perform significantly worse on recent issues from the same repositories where they excel on SWE-Bench tasks, this would strongly suggest that their success stems from memorizing specific historical issues, instance-specific memorization, rather than repository-bias memorization.
-3.2.3 External-Repo Tasks
-To further test our contamination hypothesis, we constructed tasks from repositories outside the SWE-Bench dataset but with high likelihood of training data inclusion.
-We collected the earliest available issues and pull requests from two popular repositories: pandas-dev/pandas  [^4] https://github.com/pandas-dev/pandas (123 tasks) and pytorch/pytorch  [^5] https://github.com/pytorch/pytorch (32 tasks).
-These repositories are widely used, extensively documented, and have been publicly available for years, making them prime candidates for inclusion in large-scale training datasets.
-This setting helps us recognize repository-bias memorization.
-If there is strong performance on these extra-repo tasks, this would counter the possibility that models are memorizing the instances on SWE-Bench Verified.
-However, if models to perform significantly better on SWE-Bench tasks than on these equally-exposed repository tasks, this would indicate that that models are biased toward tasks sourced from repositories included in SWE-Bench, rather than developing general coding capabilities.
+To more systematically evaluate our benchmark contamination hypothesis, we design a set of complementary test conditions for the SWE-Bench-Verified subset.
+We consider three extensions: (1) Full-SWE-Bench: randomly sampled tasks from the SWE-Bench dataset, (2) SWE-Bench Extra: tasks collected from recently created issues in repos that are included in SWE-Bench, and (3) Outside-Repo Tasks: tasks from popular repos not present in SWE-Bench.
+These additional settings enable us to assess both benchmark-specific and temporal memorization patterns.
+Full-SWE-Bench
+We randomly sample 200 instances from the full SWE-Bench (Jimenez et al.
+2024) dataset, not included in the Verified subset.
+Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these other instances, it may indicate that models have been optimized specifically for the curated subset or have seen similar patterns during training, i.e., instance-specific memorization.
+SWE-Bench Extra
+As a complementary experiment, we also collected the most recent issues from SWE-Bench repos that were mostly created after the original SWE-Bench dataset cutoff date.
+This includes 217 tasks from all repositories.
+These fresh tasks from the same repositories in SWE-Bench provide a critical test: if models perform significantly worse on recent issues from the same repos, where they excel on SWE-Bench tasks, it would strongly suggest that their success stems from instance-specific memorization, rather than repository-bias memorization.
+Outside-Repo Tasks
+To further test our hypothesis, we constructed tasks from repositories outside the SWE-Bench dataset but with a high likelihood of training data inclusion.
+We collected 245 SWE-Bench style instances from 7 popular repos: jupyter/notebook, celery/celery/, aio-libs/aiohttp, scipy/scipy, numpy/numpy, pytorch/pytorch, and pandas-dev/pandas.
+These repos are widely used, well-documented, and have been publicly available for years.
+If there is strong performance on these extra-repo tasks, this would counter the possibility that models memorize the instances on SWE-Bench Verified.
+However, if models perform significantly better on SWE-Bench tasks than on these equally-exposed repo tasks, this would indicate that models are biased toward tasks sourced from repositories included in SWE-Bench, rather than developing general coding capabilities.
 3.3 Metrics
-In this section, we details our choice of metrics used for our experiments.
+In this section, we discuss metrics used in our experiments.
 Accuracy (Acc.)
@@
 To address concerns that models might succeed through superficial pattern matching rather than genuine understanding, we introduce filtered accuracy as a control metric.
-This measures model performance exclusively on issue instances where problem descriptions do not explicitly contain file paths or import statements, identified through simple heuristics.
+This measures performance exclusively on instances where problem descriptions do not explicitly contain file paths or import statements, identified through simple heuristics.
 By removing cases where path-like strings appear in issue descriptions, filtered accuracy isolates performance that depends purely on repository knowledge and problem-solving reasoning rather than surface-level string matching.
@@
 […]
-5-gram Consecutive Accuracy (5-gram Acc.)
-To quantify the degree of verbatim memorization in model outputs, we compute 5-gram consecutive accuracy, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the ground truth.
-This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching to prevent inflation from repeated phrases.
-Given that random 5-gram matches are statistically improbable, high scores on this metric provide strong evidence of memorization rather than genuine code generation.
-This allows us to distinguish models that have memorized data patterns in the training samples.
+5-gram Consecutive Overlap (5-gram Overlap.)
+To quantify the degree of verbatim memorization in model outputs, we compute 5-gram overlap ratio, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the GT code.
+This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching that respects the occurrence count of each n-gram in the ground truth to prevent inflation from repeated phrases.
+This metric captures the extent to which models regurgitate exact token sequences.
+When models show high overlap on some benchmarks but not others, this differential pattern can indicate memorization of specific training data rather than consistent coding ability.
+$\text{N-gram Overlap (\%)}=\frac{\text{Number of matched n-grams}}{\text{Total generated n-grams}}\times 100$ (2)
+Instance-Level Verbatim Match Percentage
+For the prefix completion task, an instance is marked as compromised if the model generates at least one code hunk that exactly matches the corresponding GT implementation.
+A high instance-level verbatim match percentage suggests that many test instances may be untrustworthy, as portions of the output are memorized by the model.
+We report the percentage of compromised instances across the entire benchmark.
 3.4 Models
 We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task.
-From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), the reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor OpenAI (2025a).
-From Anthropic, our suite comprised the Claude family: Claude 3.5 Sonnet (June 20, 2024), Claude 3.7 Sonnet (February 24, 2025), Claude 4.0 Sonnet, and Claude 4.0 Opus, both officially released on May 22, 2025.
+From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor (OpenAI 2025a).
+From Anthropic, our suite comprised the Claude family: 3.5 Sonnet (June 20, 2024), 3.7 Sonnet (February 24, 2025), 4.0 Sonnet, and 4.0 Opus, both officially released on May 22, 2025.
 For consistency, OpenAI’s chat models and Claude models were configured with a maximum generation length of 2,048 completion tokens.
-For the reasoning-centric “o” family from OpenAI (i.e., o3, o3‑mini, o4‑mini), we increased the limit to 4,096 tokens.
+For the reasoning-centric “o” family (i.e., o3, o3‑mini, o4‑mini), we increased the limit to 4,096 tokens.
 All experiments used the respective provider APIs with default sampling settings to reflect typical usage.
@@
 We analyze the evidence for each pattern below.
-4.1.1 Instance-Specific memorization
-The first evidence pattern emerges from performance differences within the SWE-Bench repository ecosystem.
+Instance-Specific memorization
+The first evidence pattern emerges from performance differences within the SWE-Bench repo ecosystem.
 Models consistently achieve their highest accuracy on the curated SWE-Bench-Verified subset (60-76%), with a noticeable decline on the broader Full-SWE-Bench set (57-71%) and a further drop on newly collected SWE-Repo tasks (50-68%).
 This graduated performance decay suggests that models’ knowledge is not uniform even across the same codebases.
-The superior performance on the ”Verified” set, which is the most commonly used for evaluation and reporting, points toward instance-specific memorization, where models have been disproportionately exposed to or optimized for these specific, widely-circulated benchmark problems.
-4.1.2 Repository-Bias memorization
+The superior performance on the “Verified” set, which is the most commonly used for evaluation and reporting, points toward instance-specific memorization, where models have been disproportionately exposed to or optimized for these specific, widely-circulated benchmark problems.
+Repository-Bias memorization
 The second, more dramatic pattern appears when comparing performance between SWE-Bench and external benchmarks.
-While this performance drop is partly compounded by confounding variables, such as the significantly shorter issue descriptions in RefactorBench (Table 1), the data still points overwhelmingly to repository-level bias.
+While this performance drop is partly compounded by confounding variables, such as the significantly shorter issue descriptions in RefactorBench (Table 1), the data still points overwhelmingly to repo-level bias.
 The most compelling evidence comes from the Outside-Repo Tasks.
-These tasks were sourced from popular repositories (pandas, PyTorch) that are almost certainly well-represented in training data.
+These tasks were sourced from popular repositories that tend to be well-represented in training data.
 If models developed a general capability for popular Python projects, we would expect strong performance here.
-Instead, accuracy is uniformly lower (below 53%) than those in SWE-Bench Verified.
-The stark contrast between high performance on the SWE-Bench repositories and poor performance on these other popular repositories refutes the idea that models are simply good at ”popular projects.” It strongly supports our hypothesis of repository-bias memorization, where models have overfit to the specific architectural patterns and problem distributions of the dozen repositories in the SWE-Bench collection, failing to develop transferable skills.
-4.1.3 Model-Specific Contamination Patterns
-The results reveal differences in contamination patterns across model families and vendors:
-OpenAI chat Models: OpenAI’s chat model variants show inconsistent performance patterns across benchmarks, with no stable ranking order.
+Instead, accuracy is uniformly lower (<<53%) than that in SWE-Bench Verified.
+This stark contrast between high performance on SWE-Bench repositories and poor performance on equally popular repositories reveals that model success depends on more than just repository popularity.
+It thus supports our hypothesis of repository-bias memorization, where models have overfit to the specific architectural patterns and problem distributions of the twelve repositories in the SWE-Bench collection, failing to develop transferable skills.
+Model-Specific Contamination Patterns
+The results reveal differences in contamination patterns across model families/vendors:
+OpenAI chat Models: OpenAI’s chat models show inconsistent performance patterns across benchmarks, with no stable ranking order.
 The performance ranking of these models varies significantly between different benchmark types.
-The only consistent pattern is GPT-4o-2024-08-06, which maintains mediocre performance across most evaluations.
-OpenAI reasoning Models: While o3 demonstrates higher accuracy than o3-mini and o4-mini due to its enhanced capabilities, the performance patterns of o3-mini and o4-mini follow similar trends to the chat models.
-We cannot identify a consistently stronger model within this family for the path prediction task across different benchmarks.
-Anthropic Models: Claude models exhibit more consistent performance patterns, with newer model generations generally showing higher path prediction accuracy.
-However, Claude Sonnet 4 shows notably lower performance than other Claude family models on RefactorBench and Extra-Repo tasks, which represents an exception to the otherwise consistent generational improvement pattern.
-Overall, Claude Sonnet 4 and Opus 4 maintain relatively higher performance across most benchmarks compared to other Claude models.
-Cross-Vendor Consistency: A key finding from our evaluation of ten different LLMs is the remarkable consistency of these patterns across different model families and vendors.
-The fact that all models, regardless of vendor, exhibit the same performance hierarchy (SWE-bench Verified >>> Full-SWE-Bench >>> SWE-Repo Tasks >>> SWE-bench C#, RefactorBench, and Outside-Repo Tasks) indicates the performance disparities reflect systematic exposure patterns in training data rather than isolated vendor-specific issues.
-4.1.4 Implications for Benchmark Validity
+The only consistent pattern is GPT-4o-2024-08-06, which shows mediocre performance across most experiments.
+OpenAI reasoning Models: While o3 demonstrates higher accuracy than o3-mini and o4-mini due to its enhanced capabilities or higher degree of overfitting, the performance patterns of o3-mini and o4-mini follow similar and inconsistent trends to the chat models.
+Anthropic Models: Claude models exhibit more consistent performance patterns, with newer model generations showing higher path prediction accuracy.
+However, Claude Sonnet 4 shows notably lower performance than other Claude family models on RefactorBench, representing an exception to the otherwise consistent generational improvement pattern.
+Overall, Sonnet 4 and Opus 4 maintain relatively higher performance across most benchmarks compared to other Claude models.
+Cross-Vendor Consistency: A key finding from our evaluation of ten different LLMs is a remarkable consistency of these patterns across different model families and vendors.
+The fact that all models, regardless of vendor, exhibit the same performance hierarchy (SWE-bench Verified >> Full-SWE-Bench >> SWE-Repo Tasks >> SWE-bench C#, RefactorBench, and Outside-Repo Tasks) indicates the performance disparities reflect systematic exposure patterns in training data rather than issues isolated to a vendor.
+Implications for Benchmark Validity
 The results provide strong empirical evidence for the two forms of memorization hypothesized in our introduction.
-The performance degradation within the SWE-Bench ecosystem points to instance-specific memorization, while the more significant performance collapse on external repositories indicates repository-bias memorization.
-The dual presence of these phenomena suggests that high scores on SWE-Bench are not purely indicative of generalizable coding ability.
-Instead, they are likely inflated by these confounding factors, reflecting a specialized knowledge of the benchmark’s specific instances and repository patterns rather than transferable software engineering skills.
+The performance degradation within the SWE-Bench style benchmarks points to instance-specific memorization, while the more significant performance degradation on external repositories indicates repository-bias memorization.
+The presence of these two types of memorization suggests that high scores on SWE-Bench are not purely indicative of generalizable coding abilities for these models.
+Instead, they are likely inflated by these confounding factors.
+This shows that models possess specialized knowledge of specific tasks in SWE-Bench Verified and repository patterns, rather than or possibly in addition to transferable software engineering skills.
 Dataset Mentioned # (%) Unmentioned # (%) Total #
@@
 Full-SWE-Bench 42 (21.0%) 158 (79.0%) 200
-SWE-Repo Tasks 100 (29.1%) 244 (70.9%) 344
-SWE-bench C# 11 (14.7%) 64 (83.1%) 77
+SWE-Repo Tasks 57 (26.3%) 160 (73.7%) 217
+Outside-Repo Tasks 63 (25.7%) 182 (74.3%) 245
 RefactorBench 1 (2.6%) 38 (97.4%) 39
-Outside-Repo Tasks 26 (16.5%) 132 (83.5%) 158
-Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contains ground truth path information and those do not.
+SWE-bench C# 11 (14.7%) 64 (83.1%) 75
+Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contain ground truth path information and those that don’t.
 Figure 4: Comparison of models’ filtered accuracy on the file path identification task across benchmarks.
@@
 The results, presented in Figure 4, remove superficial string matching as a confounding variable and reveal that the underlying memorization patterns persist.
-First, models still achieve their highest accuracy on SWE-Bench Verified instances, followed by Full SWE-Bench and SWE-Repo tasks confirming a pattern of instance-specific memorization that is not dependent on clues from the prompt.
-Second, the sharp performance drop on external repositories also persists, providing definitive evidence for repository-bias memorization.
+First, models still achieve the highest accuracy on SWE-Bench Verified instances, compared to Full SWE-Bench and SWE-Repo tasks, showing a pattern of instance-specific memorization inside SWE-Bench repositories.
+Second, the sharp performance drop on external repos also persists, providing evidence for repository-bias memorization.
 The persistence of these patterns in the filtered data is our most critical finding.
 It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns.
-This suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
-Figure 6: Comparison of models’ consecutive 5-gram accuracy on function reproduction across benchmarks.
-Red ticks indicate improved verbatim similarity with the ground truth code, while blue ticks indicate a decrease.
+This suggests that current benchmark scores are inflated by these memorization effects and don’t purely reflect a model’s transferable software engineering skills.
+Model Match (%)
+Claude 3.5 Sonnet 12.1
+Claude 3.7 Sonnet 12.3
+Claude 4 Sonnet 21.4
+Claude 4 Opus 31.6
+GPT-4.1 17.4
+GPT-4o 24-05-13 18.4
+GPT-4o 24-08-16 18.2
+OpenAI o3-mini 11.7
+Table 3: Instance-level verbatim match percentage across models on SWE-Bench Verified.
+Figure 5: Comparison of models’ 5-gram overlap ratio against the ground truth code snippet on the function reproduction task across benchmarks.
 4.3 Function Reproduction Result
 We analyze model performance on the function reproduction task in this subsection.
-As shown in Figure 6, models achieve up to 35.0% consecutive 5-gram accuracy reproducing ground truth functions (the lower plot) from SWE-Bench Verified compared to up to 18.2% on instances outside the SWE-Bench ecosystem666 Due to limited resources, we only conducted experiments for o3-mini for the OpenAI reasoning models.
+As shown in Figure 5, models achieve up to 34.9% 5-gram overlap ratio reproducing ground truth functions from SWE-Bench Verified compared to up to 18.2% on instances outside the SWE-Bench ecosystem111 Due to limited resources, we were only able to conduct experiments for o3-mini from among the OpenAI reasoning models.
 .
 With limited context provided, only function names and issue descriptions, models lack sufficient information to systematically derive correct implementations.
-High consecutive 5-gram accuracy under these constraints requires reproducing exact textual sequences, indicating reliance on memorized training data rather than reasoning.
+High consecutive 5-gram overlap under these constraints requires reproducing exact textual sequences, indicating reliance on memorized training data rather than reasoning.
 Instance-Specific Memorization
-: Performance reveals a clear hierarchy that mirrors benchmark curation processes: SWE Verified (35.0%) >>> SWE Full (28.9%) >>> SWE Extra (18.3%).
-Critically, SWE Extra performance drops to the same level as external benchmarks (RefactorBench: 18.2%, Outside Tasks: 17.7%), despite originating from identical repositories as SWE Verified.
-This pattern indicates that models have specifically memorized the curated, canonical solutions that human evaluators deemed high-quality, rather than developing general familiarity with repository codebases.
-These results demonstrate that SWE-Bench Verified performance substantially reflects memorization of training sequences rather than coding competence.
-The inability to reproduce functions from less-curated instances within the same repositories, combined with performance equivalence between SWE Extra and external benchmarks, reveals that apparent programming ability is largely an artifact of exposure to the specific, high-quality solutions selected for benchmark inclusion.
+: Performance reveals a clear hierarchy mirroring benchmark curation processes: SWE Verified (34.9%) >> SWE Full (28.7%) >> SWE Extra (18.2%).
+Critically, SWE Extra performance drops to the same level as external benchmarks (RefactorBench: 18.1%, Outside Tasks: 13.9%), despite originating from identical repositories as Verified.
+This pattern indicates that models have specifically memorized the curated, canonical solutions that human evaluators deemed high-quality, rather than developing general familiarity with repositories.
+These results demonstrate that SWE-Bench Verified performance substantially reflects memorization of training sequences.
+The inability to reproduce functions from less-curated instances within the same repositories, combined with performance equivalence between SWE Extra and external benchmarks, reveals that apparent programming ability might result from an artifact of exposure to the specific, high-quality solutions selected for benchmark inclusion.
+4.4 Direct Memorization on SWE-Bench Verified
+Table 3 presents the results of our prefix completion experiment across eight SoTA language models, showing substantial evidence of training data contamination in SWE-Bench Verified, with the instance-level verbatim match percentage ranging from 11.7% to 31.6% across evaluated models.
+The Claude family exhibits a monotonic increase in memorization rates corresponding to model generation, rising from 12.1% to 31.6%.
+This progression suggests a strong correlation between model capacity and memorization.
+Notably, Claude 4 Opus demonstrates verbatim generation for nearly a third of all test instances.
+In contrast, the GPT family displays relatively stable memorization rates (17.4%-18.4%), indicating potential differences in training data curation or architectural choices that affect memorization.
+OpenAI o3-mini exhibits the lowest repro rate at 11.7%.
+These results have significant implications for benchmark validity.
+The ability of models to reproduce exact code sequences when provided merely with contextual prefixes, without any problem description or bug identification, strongly suggests that performance on these instances reflects memorization rather than algorithmic reasoning or program comprehension.
+This finding strengthens our hypothesis that a substantial portion of reported performance gains on SWE-Bench may be attributed to memorization rather than genuine advances in programming capabilities.
 5 Related Work
 5.1 Coding Benchmarks and Agents
-Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’ coding abilities, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval Jimenez et al.
-(2024); Zhuo et al.
-(2024); Xia et al.
-(2024).
-Among them, SWE-Bench has emerged as one of the most popular evaluation method for LLM’s coding abilities (Jimenez et al., 2024).
+Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval (Jimenez et al.
+2024; Zhuo et al.
+2024; Xia, Deng, and Zhang 2024).
+Among them, SWE-Bench has emerged as the most popular evaluation method for LLM’s coding abilities (Jimenez et al.
+2024).
 The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting.
 However, the public nature of GitHub data raises concerns about potential exposure during training.
-Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al., 2024; Wang et al., 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
+Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al.
+2024; Wang et al.
+2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
 5.2 Benchmark Contamination
-The issue of benchmark contamination, where models are inadvertently trained on evaluation data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding Zhou et al.
-(2023).
-To specifically combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
-Task mutation approaches aim to detect memorization by altering existing benchmark problems.
-For example, EvoEval Xia et al.
-(2024) uses LLMs to create semantic-altering transformations of HumanEval Chen et al.
-(2021) problems, while other methods Chen et al.
-(2025) use AST-based mutations to measure how performance changes with structural code modifications.
-Similarly, TaskEval Tambon et al.
-(2025) measures prompt sensitivity by creating numerous variations of the same task.
+The issue of benchmark contamination, where models are inadvertently trained on eval data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding (Zhou et al.
+2023).
+To combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
+Task mutation approaches aim to detect memorization by altering existing benchmarks.
+For example, EvoEval (Xia, Deng, and Zhang 2024) uses LLMs to create semantic-altering transformations of HumanEval (Chen et al.
+2021) problems, while other methods (Chen et al.
+2025) use AST-level mutations to measure how performance changes with structural code changes.
+Similarly, TaskEval (Tambon et al.
+2025) measures prompt sensitivity by creating variations of a task.
 These methods are effective at revealing overfitting on specific problem formulations.
-Metric-based probing approaches, in contrast, analyze the model’s outputs without changing the task itself.
-These techniques Li (2023); Ramos et al.
-(2025) include using perplexity analysis or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if a model has likely seen a specific bug or solution before.
-However, perplexity and NLL approaches cannot be applied to commercial models, for we do not have access to the model’s hidden states.
-N-gram similarity, while applicable, can be a noisy indicator for complex code patches; any functionally correct solution will inevitably share significant textual overlap with the ground truth, making it difficult to distinguish genuine problem-solving from verbatim memorization.
-While existing mutation and metric-based methods are effective for single-function code generation, they face limitations when applied to complex, multi-file software engineering benchmarks like SWE-Bench.
-Task mutation approaches require generating semantically equivalent variations of real-world GitHub issues, a significantly more challenging task than mutating isolated coding problems.
-Metric-based approaches using perplexity or NLL cannot be applied to commercial models due to lack of access to hidden states, while n-gram similarity becomes noisy for complex patches where any functionally correct solution will share substantial textual overlap with ground truth.
-Our work introduces a novel differential testing framework that addresses these limitations by using comparative performance analysis across related benchmarks as memorization proxies.
-Rather than requiring ground truth knowledge of training data exposure, we systematically compare model performance across benchmark variants and external evaluations to detect suspicious performance patterns.
-This differential testing methodology provides a systematic, reusable framework for detecting memorization in any coding benchmark without requiring access to training data or model internals.
+Metric-based probing approaches, in contrast, analyze model outputs without changing the task itself.
+These techniques (Li 2023; Ramos et al.
+2025) include using perplexity or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if model has likely seen a specific bug or solution before.
+However, these measures cannot be applied to commercial models as we do not have access to the model’s hidden states.
+N-gram similarity, while applicable, is a noisy indicator for complex code patches; any functionally correct solution will inevitably share text overlap with ground truth, making it hard to distinguish genuine problem-solving from verbatim memorization.
+Our work introduces a novel cross benchmark analysis framework that addresses these limitations by comparing models across related benchmarks as memorization proxies.
+Rather than requiring knowledge of training data exposure, we systematically compare model performance across benchmark variants and external evaluations to detect suspicious performance patterns.
+This provides a systematic, reusable framework for detecting memorization in any coding benchmark without requiring access to training data or model internals.
 6 Conclusion
-This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization.
-Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem (in both file-path identification task and function reproduction task), and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points for file-path identification task) despite their popularity and likely training data inclusion.
-The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
+This work presents evidence that current evaluations of LLM coding abilities may be compromised by benchmark contamination and memorization.
+Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific, evidenced by graduated performance decay within SWE-Bench ecosystem (in both file-path identification and function reproduction tasks), and repository-bias memorization, demonstrated by substantial performance drops on external repos (up to 47 percentage points for file-path identification) despite their popularity.
+The consistency of these patterns across different model families and vendors indicates systematic exposure patterns of SWE-Bench Verified’s data.
 These findings suggest that the improvements in SWE-Bench Verified performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities.
-The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills.
-Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than dataset-specific artifacts.
-Appendix A Appendix
-A.1 Full Prompt Template
+The field urgently needs more robust evaluation frameworks with temporal controls to prevent training data contamination, cross-repository validation to test generalization of these models beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills.
+Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than overfitting on dataset-specific artifacts.
+Appendix A SWE-Bench C# Details
+We show the repository distribution of SWE-Bench C# in Figure 6.
+This benchmark is useful for tuning agents problem-solving approach for bugfixing, feature addition, and refactoring in the C# and .NET ecosystem.
+Figure 6: Distribution of repositories in SWE-Bench C#.
+Appendix B Prompt Templates and Examples
+B.1 File-Path Identification Prompt Template
+We present the prompt template for the File-Path Identification task in Figure 7.
+The prompt begins with the repository name, followed by the issue description, which serves as the problem statement for SWE-Bench Verified tasks, and concludes with instructions for the model to reason through the problem and output a single predicted buggy file path.
+B.2 SWE-Bench Verified Task Example
+Figure 8 is a reduced example for SWE-Bench Verified instance with ID: scikit-learn__scikit-learn-10297.
+In practice, problem statements similar to this will be provided to LLMs in the file-path identification task, and the function reproduction task as a part of the prompt.
 def get_file_path_prompt(entry):
@@
 Figure 7: The complete implementation of the file path identification prompt used in our experiments.
-RidgeClassifierCV store_cv_values Issue Error: TypeError: __init__() got an unexpected keyword argument ’store_cv_values’ Repro: from sklearn import linear_model import numpy as np x = np.random.randn(100, 30) y = np.random.normal(size=100) model = linear_model.RidgeClassifierCV( alphas=np.arange(0.1, 1000, 0.1), normalize=True, store_cv_values=True ).fit(x, y) Note: store_cv_values is not a valid parameter, yet some attributes depend on it.
+RidgeClassifierCV store_cv_values Issue Error: TypeError: __init__() got an unexpected keyword argument ’store_cv_values’ Repro: 1from sklearn import linear_model 2import numpy as np 3 4x = np.random.randn(100, 30) 5y = np.random.normal(size=100) 6 7model = linear_model.RidgeClassifierCV( 8 alphas=np.arange(0.1, 1000, 0.1), 9 normalize=True, 10 store_cv_values=True 11).fit(x, y) Note: store_cv_values is not a valid parameter, yet some attributes depend on it.
 Figure 8: Minimal reproduction of RidgeClassifierCV parameter issue
-A.2 Additional Results
-[Additional tables and figures supporting the main findings]
-Model Mentioned Not Mentioned Overall
-Single GT Multi GT Overall Single GT Multi GT Overall Single GT Multi GT Overall
-claude-3-5-sonnet 88.60 71.43 85.93 51.90 44.90 50.96 61.63 52.86 60.40
-claude-3-7-sonnet 92.11 85.71 91.11 60.13 57.14 59.73 68.60 65.71 68.20
-claude-sonnet-4 91.23 85.71 90.37 65.51 53.06 63.84 72.33 62.86 71.00
-claude-opus-4 92.11 95.24 92.59 67.41 61.22 66.58 73.95 71.43 73.60
-gpt-4o (May) 91.23 85.71 90.37 51.27 55.10 51.78 61.86 64.29 62.20
-gpt-4o 88.60 80.95 87.41 53.16 51.02 52.88 62.56 60.00 62.20
-gpt-4.1 85.09 71.43 82.96 56.65 55.10 56.44 64.19 60.00 63.60
-o3-mini 91.23 85.71 90.37 60.13 63.27 60.55 68.37 70.00 68.60
-o3 92.98 76.19 90.37 71.84 65.31 70.96 77.44 68.57 76.20
-o4-mini 92.98 80.95 91.11 59.49 48.98 58.08 68.37 58.57 67.00
-Table 3: Accuracy of model-predicted file paths on SWE-Bench Verified, stratified by whether any GT path is explicitly mentioned in the problem description.
-Results are further broken down into three categories: Single GT Path (tasks where the GT patch modifies only a single execution path), Multiple GT Paths (tasks involving modifications across multiple execution paths), and Overall (all tasks).
-Among the task instances, 135 mention at least one GT path, while 365 do not.
+B.3 SWE-Bench C# Example
+Figure 9 shows the problem statement for instance: ardalis__cleanarchitecture-530 in SWE-Bench C#.
+Issue: Request to Create New ToDoItems, instance: [ardalis__cleanarchitecture-530] 1Please show how you would create new ToDoItems 2<!-- Do Not Delete This!
+feature_request_template --> 3<!-- Please search existing issues to avoid creating duplicates.
+--> 4 5<!-- Describe the feature youd like.
+--> 6 7In this template, you aggregate ToDoItems within the ProjectAggregate.
+However, you also stop short of showing how you would, via API, author new ToDoItem instances.
+I think it would be beneficial to show how you would create these new child items of a project.
+8 9I think it would be wonderful to see: 10- How you would define this in an ApiEndpoint.
+11- How you would incorporate this within the ProjectAggregate.
+12- How you would codify a Many-to-Many relationship as opposed to the One-To-Many relationship between Project and ToDoItem.
+13 - Is this implied by the Project <-> Contributor relationship?
+14 - What if the Contributor to a Project had a title within that Project?
+Senior Contributor vs Junior Contributor?
+If that were the case, what ApiEndpoint would the management of that be within your example domain?
+15 16Thanks for such a great template!
+Figure 9: Example issue description requesting creation of new ToDoItem instances
+B.4 Function Reproduction Task Prompt Template
+We present the prompt template for the Function Reproduction task in Figure 10.
+The prompt begins with the repository name, followed by hints indicating which functions need to be implemented, and only the names and the file paths are included in the hints.
+It then provides the complete file content, where the buggy functions are removed and replaced by comments marking the index of each missing function.
+Finally, the prompt instructs the model to reason through the task and output the reconstructed content for all buggy functions identified in the provided file.
 You are provided with a code repository and an issue description.
@@
 Figure 10: The prompt template used to test models’ ability to reproduce ground truth function without function specification.
+Appendix C Additional Experiment Results
+C.1 Comparison of 5-Gram Overlap for Buggy V.S. Ground Truth Code Snippet
+Figure 11: Difference in models’ 5-gram overlap ratio when comparing against the ground truth code snippet versus the buggy code snippet in the function reproduction task across benchmarks.
+A higher difference indicates that the model’s solution is more similar to the ground truth.
+We also report the difference of 5 gram overlap
+$\Delta_{5}=\text{overlap}_{5}(\hat{f},f_{\text{GT}})-\text{overlap}_{5}(\hat{f},f_{\text{buggy}}),$
+where $\text{overlap}_{5}(x,y)$ is the 5 gram overlap of $x$ and $y$.
+Positive values indicate the model output more closely matches the patched implementation; negative values indicate closer similarity to the buggy code.
+Figure 11 summarizes $\Delta_{5}$ across benchmarks and model families. Three broad patterns emerge.
+(1) Provider‑level asymmetry on legacy benchmarks.
+Anthropic’s Claude models show consistently positive differentials on SWE‑Bench Verified, Full-SWE-Bench, and Outside-Repo Tasks (roughly +2 to -6 percentage points), suggesting that these models retain verbatim shards of the patched implementations.
+OpenAI models yield smaller positives (generally less than +2pp) over the same sets, indicating weaker memorization of those specific fixes.
+(2) Freshness stress test.
+All models, except Claude 3.5 Sonnet and GPT 4.1, produce negative $\Delta_{5}$ on SWE‑Bench Extra.
+Because this split is mined largely from post‑cutoff commits, the patches are unlikely to be in training data.
+The systematic shift toward the buggy reference therefore supports our contamination‑mitigation claim: when the patched code is unseen, models do not memorize it from context.
+(3) Language / ecosystem skew.
+On the SWE C# subset, OpenAI’s GPT‑4* family exhibits the largest positive differentials (~+3 to 5pp), exceeding Claude.
+This may suggest heavier exposure to patched C# / .NET code in OpenAI training data (or lighter deduplication of that ecosystem) compared to Anthropic.
+Taken together, the function reproduction experiment provides a direct, contrastive probe of code memorization: when historical (patched) code is in a model’s training mix, we observe measurable positive $\Delta_{5}$; when the code post‑dates the model (SWE‑Bench Extra), the signal disappears or reverses.
+Additionally, the high tendency towards the fixed version of code on SWE-Bench Verified and SWE-Bench C# indicates higher possibility of benchmark memorization on these two settings.
+These results underscore the necessity of freshness‑controlled benchmarks when interpreting coding model performance and memorization risk.